OpenStudy (karatechopper):

sqrt of 108 = 6 times sqrt of 3?

OpenStudy (karatechopper):

am i right didi?

OpenStudy (anonymous):

NOOOOOOOOOOOOOOOOOP

OpenStudy (karatechopper):

then what is it?

OpenStudy (diyadiya):

Yes!!

OpenStudy (karatechopper):

i am right?!

OpenStudy (anonymous):

NOOOO D:

OpenStudy (karatechopper):

prove i am wrong michu

OpenStudy (kinggeorge):

You are correct. \(\sqrt{108}=6\sqrt{3}\) Since \(6^2 \cdot 3=36 \cdot 3=108\)

OpenStudy (karatechopper):

i wanna see what i did wrong mich

OpenStudy (karatechopper):

u

OpenStudy (karatechopper):

so show me ur work

OpenStudy (diyadiya):

Just square both sides :)

OpenStudy (karatechopper):

lol didi i know how to do it

OpenStudy (diyadiya):

Good :)

OpenStudy (karatechopper):

yep:D